# Matching Free Depth Recovery from Structured Light

### v1

@@ Abstract
We present a novel approach for depth estimation from images captured by structured light systems. Unlike many previous methods that rely on image matching process, our approach uses a density voxel grid to represent scene geometry, which is trained via self-supervised differentiable volume rendering. Our method leverages color fields derived from projected patterns in structured light systems during the rendering process, enabling the isolated optimization of the geometry field. This contributes to faster convergence and high-quality output. Additionally, we incorporate normalized device coordinates (NDC), a distortion loss, and a novel surface-based color loss to enhance geometric fidelity. Experimental results demonstrate that our method outperforms existing matching-based techniques in geometric performance for few-shot scenarios, achieving approximately a 60% reduction in average estimated depth errors on synthetic scenes and about 30% on real-world captured scenes. Furthermore, our approach delivers fast training, with a speed roughly three times faster than previous matching-free methods that employ implicit representations.
1 Introduction
@@
With the advent of structured-light cameras such as Kinect V2 and Intel RealSense [17], structured light systems have become a powerful solution for range sensing in various applications [4, 31, 32, 42]. Typically, a monocular structured light system consists of one camera and one projector, both with calibrated intrinsic and extrinsic parameters. By projecting randomly or manually designed patterns into 3D space, the system extracts depth information by analyzing the deformation of these patterns in captured images. Classical algorithms in structured light aim to establish robust correspondences across multiple projected patterns, as illustrated in Fig. 1.
@@
Figure 1: An illustration of a monocular structured light system. Depth is computed by pixel correspondences between camera and projector. These correspondences are typically calculated through image matching algorithms.
@@
The 3D scanning process in structured light systems often requires balancing the trade-off between scan precision and the number of projected patterns or captured images. Increasing the number of captured images can improve the accuracy of correspondence matching. However, more projected patterns also lead to longer acquisition times, which limits the applicability of structured light systems in general scenes, especially those with undesired motion and short exposure times. To address this, researchers have developed techniques to embed richer information within a limited set of patterns [46, 12, 11, 14, 23]. These methods use sophisticated patterns that encode temporal or spatial features to mitigate matching uncertainties. By decoding these features from captured images, structured light systems can determine the correspondences needed for accurate depth estimation [16, 18, 29, 30].
@@
Despite these advances, designing features that can produce accurate dense depth maps while remaining resilient to environmental influences remains a significant challenge. Traditional structured light depth recovery methods rely on the accuracy of the matching algorithm between projected patterns and captured images. Any errors in the matching process, such as blurring or occlusions, can introduce substantial inaccuracies in the final depth maps. Recent deep learning techniques offer solutions by using neural networks to tackle matching uncertainties [15, 28, 26]. Most of these methods train a general model that is then used to directly predict depth. While these models can generate dense depth maps, they require extensive training datasets, which can significantly impact network performance. Constructing such datasets is challenging in structured light systems, given the diversity of devices and pattern configurations involved.
@@
Our work draws inspiration from recent success in using classic voxel grids to explicitly store scene geometries [9, 13, 47]. Instead of focusing on designing or learning robust matching features, we introduce a novel framework based on volume rendering. Specifically, our approach employs a voxel grid to represent the volume density of the target 3D scene. Through a fully differentiable rendering process, we generate images from the camera’s viewpoint, calculating color from projected patterns, and establish an straight training pipeline with a direct loss function between the captured images and their rendered counterparts. Once training converges, the volume density of the 3D space is obtained, enabling the extraction of both the scene’s geometry and the depth map via simple volumetric rendering. The rendering pipeline in our approach is similar to NeRF-based techniques used in view synthesis tasks [22, 48, 49]. While these methods achieve impressive results in image synthesis from passive views, they face limitations in geometry recovery due to the need to jointly estimate radiance and geometry fields from captured images [24, 38, 44, 45]. In structured light systems, however, the radiance field is predetermined by the projected patterns, allowing us to focus solely on optimizing the geometry field for high-quality depth estimation.
@@
We leverage constraints from the projected light field to optimize the 3D voxel grid in a monocular structured light setup. This process includes a rendering mechanism that incorporates these projected pattern constraints during ray sampling. Additionally, we introduce a distortion loss to accelerate training and a surface color loss to enhance geometric accuracy. Our experimental results demonstrate that, with as few as six randomly generated binary patterns, our method significantly outperforms existing matching-based techniques in terms of geometric accuracy. Specifically, it reduces the average estimated depth error by approximately 60% on synthetic scenes and by around 30% on real-world captured scenes. Furthermore, our approach offers a substantial advantage in training efficiency, achieving a training speed nearly three times faster than methods that rely on implicit geometry representations.
In summary, the main contributions of out work are as follows:
@@
  • We propose a novel matching free framework for depth estimation in structured light systems that eliminates the need for extensive training datasets or specifically designed patterns, which are typically required in traditional matching-based techniques.
@@
  • Our method incorporates color information from the projected patterns as existing color field to train the voxel grid, which facilitates faster convergence and improves geometry estimation performance.
  • By using voxel grids for geometry representation in our training process, we achieve efficient volume rendering, which accelerates the overall training optimization process.
2 Related Work
@@
Temporal-encoding structured light systems. Temporal-encoding patterns are widely used in structured light systems for static scene reconstruction, as they enable unique decoding at each camera pixel. The plane of light is in a unique location at each time instant, and can be used to recover depth. A number of strategies for coding space with a time sequence of light patterns have been proposed, including binary strips[33], gray codes [25, 1, 40], grid patterns [19], fringe patterns [16, 18, 30, 37], and phase measurements [39, 51]. Modern structured light systems achieve high accuracy and resilience to noise when sufficient patterns are projected. However, this accuracy declines significantly in few-shot scenarios where only a limited number of patterns can be projected. To address this limitation, researchers have explored enhanced coding strategies, such as using specialized devices [1, 36, 40], complex pattern designs [19, 33], or global optimization methods during post-processing [16, 18, 29, 30]. Despite these efforts, manually modeling uncertain factors often introduces unstable outliers in the depth map estimation.
@@
Learning-based matching in structured light systems. Learning-based methods have been increasingly applied to address the matching problem in structured light systems, as demonstrated in studies like  [7, 8]. Some other works employ deep networks to directly predict disparity from image-pattern pairs within these systems[15, 28, 50]. However, these approaches require extensive training datasets, and the scarcity of publicly available benchmarks tailored to the unique characteristics of structured light—such as specific pattern designs and parameter configurations—presents a significant challenge. Besides, these learning-based method often encounter domain shift issues when applied to real scene settings.
@@
Voxel grid representation for 3D geometry. Voxel grids model 3D objects by dividing space into a regular array of volumetric units, or voxels, each storing data like color, density, or material properties. This approach is advantageous for capturing complex geometries and internal structures, especially when surface-based methods, such as meshes, fall short. Noteworthy applications of voxel grids in deep learning include VoxNet by Maturana and Scherer [21], a 3D CNN that operates on voxel grids for object recognition. Further advancing this integration, VoxGRAF by Schwarz et al. [34] introduces a sparse voxel grid framework for 3D-aware image synthesis, efficiently rendering views with 3D consistency and visual fidelity by combining sparse grids with 3D convolutions, progressive growing, and free-space pruning. DVGO by Cheng Sun et al. [35] apply voxel grids to model 3D geometry in neural rendering pipeline [22], which achieves NeRF-comparable quality and converges rapidly from scratch. In this work, we also adopt voxel grid structure to develop a tailored training framework for high-quality depth reconstruction within structured light systems.
3 Methods
Figure 2: Our pipeline for matching free depth recovery in structured light systems. Sampled rays are rendered through volume rendering process using voxel grid density. Each sampled point on the ray is converted into projector space to query the color. The ray’s rendered color and the color of its estimated surface point are both compared with the captured image to train our voxel grid.
@@ 3 Methods
Our approach focuses on monocular structured light systems, consisting of a single camera and a projector to capture images for depth estimation. Given a set of patterns $\left\{\mathbf{P}_{i}\right\}_{i=1,2,\ldots N}$, the projector sequentially projects them onto the scene, while the camera captures images $\left\{\mathbf{I}_{i}\right\}_{i=1,2,\ldots N}$. Here, $N$ represents the number of projected patterns. With the known intrinsic and extrinsic parameters of both the camera and the projector, our goal is to reconstruct the depth map $\mathbf{D}$ from the camera viewpoint. As it is under monocular setting, we can simply set the extrinsic matrix of the camera to identity matrix, i.e. $c2w=\mathbf{I}_{4\times 4}$.
@@
Traditional matching-based methods attempt to define a function which use estimated point-by-point correspondences to directly compute the depth map. In contrast, we introduce a novel matching free framework, as shown in Fig. 2. We first build a density voxel grid to store geometric information of the input scene (Sec. 3.1). This is followed by a differentiable volume rendering process to generate camera-observed images (Sec. 3.2). Finally, the network is trained using several loss functions to compare the rendered images with the captured ones (Sec. 3.3).
3.1 Density Voxel Grid
@@
where $x$ denotes the queried 3D point, $V$ represents the voxel grid, $C$ is the modality dimension, and $N_{x},N_{y},N_{z}$ is total number of voxels on each dimension. We use trilinear interpolation in our method.
Density voxel grid $\boldsymbol{V}^{\text{(density) }}$ is a special case with $C=1$, which stores the density values for volume rendering (see Eq.(4)). Similar to DVGO [35], we use $\ddot{\sigma}\in\mathbb{R}$ to denote the raw density stored by the voxel grid. We then use the shifted softplus from Mip-NeRF [2] to apply the density activation:
@@
All grid values in $\boldsymbol{V}^{\text{(density) }}$ are initially set to zero. To ensure that all sampled points on rays are visible to the camera at the start, we initialize the accumulated transmittance $T_{i}\approx 1$ by setting the shift $b$ to
@@
where $\alpha^{(\text{init })}$ serves as a hyperparameter, and $\delta$ is the step size. Employing the softplus function instead of ReLU is essential for optimizing voxel density, as ReLU may irreversibly set it to zero when a voxel value is falsely set to negative.
3.2 Rendering with Projected Patterns
For one of the captured image $\mathbf{I}_{j}$, we define its corresponding rendered image as $\widehat{\mathbf{I}}_{j}$. To render the color of a pixel $\widehat{\mathbf{I}}_{j}(\boldsymbol{r})$, we cast the ray $\boldsymbol{r}$ from the camera center through the pixel. $K$ points are then sampled on $\boldsymbol{r}$ between the predefined near and far planes. The $K$ ordered sampled points are then used to query for their densities and colors $\left\{\left(\sigma_{i},\boldsymbol{c}_{i}\right)\right\}_{i=1}^{K}$. Finally, the $K$ queried results are accumulated into a single color with the volume rendering equation [22]:
@@ 3.2 Rendering with Projected Patterns
We use post-activation strategy from [35] to calculate the density of the sampled point $\boldsymbol{x}_{i}$. That means we first use trilinear interpolation to get the raw density value:
@@
We then employ the softplus function (Eq.(2)) to get $\sigma_{i}$, and finally use it to calculate rendered results. DVGO [35] has shown that post-activation strategy can produce a sharp linear boundary.
In our approach, the color of the sampled point $\boldsymbol{x}_{i}$ is not estimated from voxel grid or network architecture. Instead, we utilize prior knowledge from our projected patterns to determine the point color using a reprojection function $\pi$. Specifically, the color $\boldsymbol{c}_{i}$ can be directly calculated from the projected pattern $\mathbf{P}_{j}$ through:
@@
3.3 Loss Functions
Figure 3: Visualization of our photometric loss $\mathcal{L}_{\text{photo }}$, distortion loss $\mathcal{L}_{\text{dist }}$, and surface color loss $\mathcal{L}_{\text{surface}}$. The photometric loss encourages accurate color in rendered images, while both distortion loss and surface color loss enforces a singular peak along the ray.
During the training process, we pick a batch of $M$ pixels from each captured image, and sample $K$ points on each corresponding ray. We then calculate the color of these rays to generate the rendered color $\widehat{\mathbf{I}}_{j}(\boldsymbol{r})$ for each image $j$. We first minimize a per-pixel loss function that quantifies the difference between $\widehat{\mathbf{I}}(\boldsymbol{r})$ and the color $\mathbf{I}(\boldsymbol{r})$ from captured image. The photometric MSE is defined as

### v2

# Matching-Free Depth Recovery from Structured Light

@@ Abstract
We introduce a novel approach for depth estimation using images obtained from monocular structured light systems. In contrast to many existing methods that depend on image matching, our technique employs a density voxel grid to represent scene geometry. This grid is trained through self-supervised differentiable volume rendering. Our method leverages color fields derived from the projected patterns in structured light systems during the rendering process, facilitating the isolated optimization of the geometry field. This innovative approach leads to faster convergence and high-quality results. Additionally, we integrate normalized device coordinates (NDC), a distortion loss, and a distinctive surface-based color loss to enhance geometric fidelity. Experimental results demonstrate that our method outperforms current matching-based techniques in terms of geometric performance in few-shot scenarios, achieving an approximately 30% reduction in average estimated depth errors for both synthetic scenes and real-world captured scenes. Moreover, our approach allows for rapid training, being approximately three times faster than previous matching-free methods that utilize implicit representations.
Index Terms:
Structured Light, Depth Reconstruction, Volume Rendering, Voxel Grid.
I Introduction
@@
The acquisition of precise depth measurements constitutes a core technical challenge in modern perception pipelines. With the advent of structured-light cameras such as Kinect V2 and Intel RealSense [1], structured light systems have become a powerful solution for depth sensing in various applications [2, 3, 4]. Typically, a monocular structured light system consists of one camera and one projector with calibrated intrinsic and extrinsic parameters. By projecting randomly or manually designed patterns into 3D space, the system extracts depth information by analyzing the deformation of these patterns in captured images. Classical algorithms in structured light aim to establish robust correspondences across multiple projected patterns. We illustrate this monocular structured light system in Fig. 1(a). The disparity map is computed between the captured image and the projected pattern. Since disparity is inversely proportional to depth, the depth map can be recovered through triangulation, given the known geometric calibration between the camera and the projector. Both disparity maps and depth maps encode scene geometry and are mathematically convertible. Therefore, in this paper, we treat disparity and depth interchangeably without explicit distinction.
((a))
((b))
@@
Figure 1: (a) An illustration of a classic monocular structured light system. Depth is computed by pixel correspondences between the camera and the projector. These correspondences are typically calculated through image matching algorithms. (b) Our monocular structured light system. During the rendering process, we sample points along each camera ray and project them onto the pattern plane to retrieve their corresponding color values.
@@
The 3D scanning process in structured light systems often requires balancing the trade-off between scan precision and the number of projected patterns or captured images. Increasing the number of captured images can improve the accuracy of correspondence matching. However, more projected patterns also lead to longer acquisition times, which limits the applicability of structured light systems in general scenes, especially those with undesired motion and short exposure times. To address this, researchers have developed techniques to embed richer information within a limited set of patterns [5, 6, 7]. These methods use sophisticated patterns that encode temporal or spatial features to mitigate matching uncertainties. By decoding these features from captured images, structured light systems can determine the correspondences needed for accurate depth estimation [8, 9, 10, 11].
@@
Despite these advances, designing features that can produce accurate dense depth maps while remaining resilient to environmental influences remains a significant challenge. Traditional structured light depth recovery methods rely on the accuracy of the matching algorithm between projected patterns and captured images. Any errors in the matching process, such as blurring or occlusions, can introduce substantial inaccuracies in the final depth maps. Recent deep learning techniques offer solutions using neural networks to tackle matching uncertainties [12, 13, 14]. Most of these methods formulate the depth estimation process with a convolutional neural network. Schreiberhuber et al. [15] formulate depth estimation as a regression problem and address it by discretizing the regression into a series of classification sub-tasks, which were solved using multi-layer perceptrons (MLPs). While these models can generate dense depth maps, they require extensive training datasets, significantly impacting network performance. Given the diversity of devices and pattern configurations, constructing such datasets is challenging in structured light systems.
@@
Our work draws inspiration from recent success in using classic voxel grids to explicitly store scene geometries [16, 17, 18]. Instead of focusing on designing or learning robust matching features, we introduce a novel framework based on volume rendering. Specifically, our approach employs a voxel grid to represent the volume density of the target 3D scene. Through a fully differentiable rendering process, we generate images from the camera’s viewpoint, calculating color from projected patterns (as shown in Fig. 1(b)), and establish a straight training pipeline with a direct loss function between the captured images and their rendered counterparts. Once training converges, the volume density of the 3D space is obtained, enabling the extraction of both the scene’s geometry and the depth map via simple volumetric rendering. The rendering pipeline in our approach is similar to NeRF-based techniques used in view synthesis tasks [19, 20, 21]. While these methods achieve impressive results in image synthesis from passive views, they face limitations in geometry recovery due to the need to jointly estimate radiance and geometry fields from captured images [22, 23, 24, 25]. In structured light systems, however, the radiance field is predetermined by the projected patterns, allowing us to focus solely on optimizing the geometry field for high-quality depth estimation.
@@
We leverage constraints from the projected light field to optimize the 3D voxel grid in a monocular structured light setup. This process includes a rendering mechanism that incorporates these projected pattern constraints during ray sampling. Additionally, we introduce a distortion loss to accelerate training and a surface color loss to enhance geometric accuracy. Our experimental results demonstrate that, with as few as six randomly generated binary patterns, our method significantly outperforms existing matching-based techniques in terms of geometric accuracy. Specifically, it reduces the average estimated depth error by approximately 30% on both synthetic scenes and real-world captured scenes, when compared with previous matching-based methods.
There are existing methods that utilize volume rendering to address depth recovery tasks in structured light systems. NFSL [26] utilizes a neural field to recover the scene’s geometry with a moving camera, extracting correspondences between different input camera views and projected patterns. Our task is quite different from theirs, as our method operates under the monocular structured light system with a single-camera pose. There are also methods considering monocular settings  [27, 28]. They train a signed distance function(SDF) as the geometry representation, and apply the marching cubes algorithm [29] to extract the depth map. While SDFs can effectively enforce surface smoothness through their implicit continuous formulation, they are not very suitable for this monocular task, as their watertight surfaces inherently result in two surface points along each camera ray, which does not align with the expected characteristics of depth maps (2D height fields). Additionally, employing SDFs in this task leads to difficulties in capturing fine details, especially in regions with sharp depth discontinuities like object boundaries. With only a single viewpoint available, the trained SDF represents the scene’s geometry as a single watertight surface, which tends to smooth out geometric discontinuities, making it difficult to recover depth in those areas accurately. In contrast, our explicit voxel grid representation is not constrained by such global modeling, allowing for more localized control, enabling finer precision in areas with significant depth variation. Furthermore, our approach offers a substantial advantage in training efficiency, achieving a nearly three times faster training speed than these methods that rely on implicit geometry representations. We visualize these differences between SDFs and voxel grids in Fig. 2.
Figure 2: Visualization of different geometry representations at object edges of the input scene. The depth map can be treated as a 2D height field with a single value at each 2D location represented by a camera ray. The SDFs generate watertight surfaces, resulting in two surface points along each ray, and leading to difficulties in representing sharp edges. Voxel grids model scene properties at discrete volumetric locations, offering superior flexibility in capturing sharp edges.
In summary, the main contributions of our work are as follows:
@@
  • We propose a novel matching-free framework for depth estimation in structured light systems that eliminates the need for extensive training datasets or specifically designed patterns typically required in traditional matching-based techniques.
@@
  • Our method incorporates color information from the projected patterns as an existing color field to train the voxel grid, facilitating faster convergence and improving geometry estimation performance.
  • By using voxel grids for geometry representation in our training process, we achieve efficient training speed and good geometry performance, outperforming existing rendering-based methods that utilize implicit representations.
II Related Work
@@
Temporal-encoding structured light systems. Temporal-encoding patterns are widely used in structured light systems for static scene reconstruction, as they enable unique decoding at each camera pixel. The plane of light is in a unique location at each time instant, and can be used to recover depth. A number of strategies for coding space with a time sequence of light patterns have been proposed, including binary strips[30], gray codes [31, 32, 33], grid patterns [34], fringe patterns [8, 10, 35], and phase measurements [36, 37]. Modern structured light systems achieve high accuracy and resilience to noise when sufficient patterns are projected. However, this accuracy declines significantly in few-shot scenarios where only a limited number of patterns can be projected. To address this limitation, researchers have explored enhanced coding strategies, such as using specialized devices [32, 38, 33], complex pattern designs [34, 30], or global optimization methods during post-processing [8, 39, 10]. Despite these efforts, manually modeling uncertain factors often introduces unstable outliers in the depth map estimation.
@@
Learning-based matching in structured light systems. Recently, learning-based techniques have become popular in many areas and have achieved tremendous success. Deep-learning-based approaches have also been adopted to address the matching problem in structured light systems, as demonstrated in studies like  [40, 41]. Some of these methods formulate the depth estimation process with a convolutional network. Some other works employ deep networks to directly predict disparity from image-pattern pairs within these systems[12, 13, 42]. Schreiberhuber et al. [15] formulate the depth estimation task as a regression problem. However, these approaches require extensive training datasets, and the scarcity of publicly available benchmarks tailored to the unique characteristics of structured light—such as specific pattern designs and parameter configurations—presents a significant challenge. Besides, these learning-based methods often encounter domain shift issues when applied to real scene settings.
@@
Voxel grid representation for 3D geometry. Voxel grids model 3D objects by dividing space into a regular array of volumetric units, or voxels, each storing data like color, density, or material properties. This approach is advantageous for capturing complex geometries and internal structures, especially when surface-based methods, such as meshes, fall short. Noteworthy applications of voxel grids in deep learning include VoxNet by Maturana and Scherer [43], a 3D CNN that operates on voxel grids for object recognition. Further advancing this integration, VoxGRAF by Schwarz et al. [44] introduces a sparse voxel grid framework for 3D-aware image synthesis, efficiently rendering views with 3D consistency and visual fidelity by combining sparse grids with 3D convolutions, progressive growing, and free-space pruning. DVGO by Cheng Sun et al. [45] applies voxel grids to model 3D geometry in a neural rendering pipeline [46], which achieves NeRF-comparable quality and converges rapidly from scratch. In this work, we also adopt a voxel grid structure to develop a tailored training framework for high-quality depth reconstruction within structured light systems.
III Methodology
@@
Our approach focuses on monocular structured light systems, consisting of a single camera and a projector to capture images for depth estimation. Given a set of patterns $\left\{\mathbf{P}_{i}\right\}_{i=1,2,\ldots N}$, the projector sequentially projects them onto the scene, while the camera captures images $\left\{\mathbf{I}_{i}\right\}_{i=1,2,\ldots N}$. Here, $N$ represents the number of projected patterns. With the known intrinsic and extrinsic parameters of the camera and the projector, we aim to reconstruct the depth map $\mathbf{D}$ from the camera viewpoint. As it is under monocular setting, we can simply set the extrinsic matrix of the camera to the identity matrix.
@@
Traditional matching-based methods attempt to define a function that uses estimated point-by-point correspondences to directly compute the depth map. In contrast, we introduce a novel matching-free framework, as shown in Fig. 3. We first construct a density voxel grid to store the geometric information of the input scene (see Section III-A). Next, we employ a differentiable volume rendering process using the voxel grid and the projected patterns to generate images under the camera view (see Section III-B). During the optimization process, we utilize several loss functions to compare the rendered images with the captured ones and to encourage voxel densities to be compact and sparse (see Section III-C). Finally, we introduce the NDC parameterization, which reallocates the voxel grid’s density to better align with the geometry of perspective projection (see Section III-D).
III-A Density Voxel Grid
@@
where $\boldsymbol{x}$ denotes the queried 3D point, $\boldsymbol{V}$ represents the voxel grid, $C$ is the modality dimension, and $N_{x},N_{y},N_{z}$ is total number of voxels on each dimension. We use trilinear interpolation in our method.
Density voxel grid $\boldsymbol{V}^{\text{(density) }}$ is a special case with $C=1$, which stores the density values for volume rendering (see Eq.(4)). To optimize voxel density directly, we use $\ddot{\sigma}\in\mathbb{R}$ to denote the raw density stored by the voxel grid and use the shifted $\operatorname{softplus}$ from Mip-NeRF [47] to apply the density activation:
@@
All grid values in $\boldsymbol{V}^{\text{(density) }}$ are initially set to zero. Employing the $\operatorname{softplus}$ function instead of $\operatorname{ReLU}$ is essential for optimizing voxel density, as $\operatorname{ReLU}$ may irreversibly set it to zero when a voxel value is falsely set to negative.
Figure 3: Our pipeline for matching-free depth recovery in structured light systems. Sampled rays are rendered through a volume rendering process using voxel grid density. Each sampled point along a ray is projected into the projector space to retrieve its corresponding color value from the structured light pattern. To supervise the voxel grid optimization, both the rendered color of the entire ray and the color of its estimated surface point are compared against the captured image.
To ensure that all sampled points on rays are visible to the camera at the start, we initialize the accumulated transmittance $T_{i}\approx 1$ by setting the shift $b$ to
@@
where $\alpha^{(\text{init })}$ serves as a hyperparameter, and $\delta$ is the step size.
III-B Rendering with Projected Patterns
For one of the captured images $\mathbf{I}_{j}$, we define its corresponding rendered image as $\widehat{\mathbf{I}}_{j}$. To render the color of a pixel $\widehat{\mathbf{I}}_{j}(\boldsymbol{r})$, we cast the ray $\boldsymbol{r}$ from the camera center through the pixel. $K$ points are then sampled on $\boldsymbol{r}$ between the predefined near and far planes. The $K$ ordered sampled points are then used to query for their densities and colors $\left\{\left(\sigma_{i},\boldsymbol{c}_{i}\right)\right\}_{i=1}^{K}$. Finally, the $K$ queried results are accumulated into a single color with the volume rendering equation [19]:
@@ III-B Rendering with Projected Patterns
We use the post-activation strategy from [45] to calculate the density of the sampled point $\boldsymbol{x}_{i}$. That means we first use trilinear interpolation to get the raw density value:
@@
We then employ the softplus function (Eq.(2)) to get $\sigma_{i}$, and finally use it to calculate rendered results. DVGO [45] has shown that the post-activation strategy can produce a sharp linear boundary.
In our approach, the color of the sampled point $\boldsymbol{x}_{i}$ is not estimated from the voxel grid or the training schedule. Instead, we utilize prior knowledge from our projected patterns to determine the point color using a re-projection function $\pi$. Specifically, the color $\boldsymbol{c}_{i}$ can be directly calculated from the projected pattern $\mathbf{P}_{j}$ through:
@@
III-C Loss Functions
During the training process, we pick a batch of $M$ pixels from each captured image, and sample $K$ points on each corresponding ray. We then calculate the color of these rays to generate the rendered color $\widehat{\mathbf{I}}_{j}(\boldsymbol{r})$ for each image $j$. We first minimize a per-pixel loss function that quantifies the difference between $\widehat{\mathbf{I}}(\boldsymbol{r})$ and the color $\mathbf{I}(\boldsymbol{r})$ from the captured image. The photometric MSE is defined as